### v1

# Data-Driven Estimation of Quadrotor Motor Efficiency via Residual Minimization

Abstract
A data-driven framework is proposed for online estimation of quadrotor motor efficiency via residual minimization. The problem is formulated as a constrained nonlinear optimization that minimizes trajectory residuals between measured flight data and predictions generated by a quadrotor dynamics model. A sliding-window strategy enables online estimation, and the optimization is efficiently solved using an iteratively reweighted least squares (IRLS) scheme combined with a primal-dual interior-point method, with inequality constraints enforced through a logarithmic barrier function. Robust z-score weighting is employed to reject outliers, which is particularly effective in motor clipping scenarios where the proposed estimator exhibits smaller spikes than an EKF baseline. Compared to traditional filter-based approaches, the batch-mode formulation offers greater flexibility by selectively incorporating informative data segments. This structure is well-suited for onboard implementation, particularly for applications such as fault detection and isolation (FDI), health monitoring, and predictive maintenance in aerial robotic systems. Simulation results under various degradation scenarios demonstrate the accuracy and robustness of the proposed estimator.
I Introduction
A quadrotor is a highly agile aerial robotic platform that relies on rapid motor response for stable and precise flight. The performance of its propulsion system directly impacts energy efficiency, flight time, and maneuverability. Motor efficiency, a critical yet often unmeasured factor, can deteriorate due to elevated temperatures, component aging, mechanical wear, and battery voltage fluctuations. An effective estimator holds significant potential for fault detection and isolation (FDI), early failure prediction, and maintenance alerts, ultimately mitigating in-flight failure risks.
Extensive research on system identification for quadrotors has been conducted to improve control accuracy and state estimation. Most prior works focus on identifying parameters such as mass, inertia, or center of mass, and detecting actuator faults. For example, [1] employed an IMU combined with motor speed commands and an Unscented Kalman Filter (UKF) to estimate mass and inertia using Newton-Euler dynamics, while [2] used a Kalman Filter (KF) combined with Recursive Least Squares (RLS) to estimate the moment of inertia. [3] utilized UKF for estimating inertia, velocity, and angular velocity, and [4] compared least-squares, EKF, and instrumental variable methods for mass estimation. Other approaches include [5], who proposed a robust geometric control framework for quadrotors that incorporates online parameter estimation, and [6] employed UKF and EKF to estimate geometric and inertial parameters. [7] used diffusion models to learn unmodeled effects such as aerodynamic disturbances, mass uncertainty, and sensor noise. For fault detection, methods such as neural networks [8], augmented-state Kalman filters [9], sparse identification and Thau’s observer [10], and nonlinear observers with sliding mode control [11] have been proposed. [12] further demonstrated a data-driven system identification framework using Maximum A Posteriori (MAP) estimation to jointly estimate motor delays and inertial properties. However, relatively little attention has been given to the estimation of motor efficiency for quadrotors, despite its critical impact on energy consumption, stability, and flight time. Among these studies, our previous work [13] introduced a UKF-based estimator integrated into a cooperative multirotor system.
This work proposes a batch-mode, optimization-based framework for estimating quadrotor motor efficiency by minimizing trajectory residuals. While not explored in this study, such formulations also offer potential for integration with learning-based methods. The estimation problem is posed as a least-squares optimization with inequality constraints to ensure the efficiency factors remain within a physically valid range (typically $0\leq\eta\leq 1$). To solve this constrained optimization problem, a primal-dual interior-point method is adopted, augmented with an outer IRLS loop that applies robust z-score weighting to suppress outliers. The inequality constraints are enforced through a logarithmic barrier function, and the interior-point framework was originally introduced by Karmarkar [14] and later extended to nonlinear problems using barrier methods by Fiacco and McCormick [15]. A comprehensive introduction of interior-point techniques, including log-barrier formulations, can be found in Boyd and Vandenberghe [16]. The main contributions of this work are summarized as follows:
  • $\bullet$
    A nonlinear optimization-based estimator with bounded constraints is developed to identify quadrotor motor efficiency by minimizing residuals between measured and predicted trajectories.
  • $\bullet$
    A sliding-window scheme with a primal-dual interior-point method and IRLS reweighting improves efficiency and robustness by rejecting outliers and reducing spikes.
  • $\bullet$
    The proposed approach is validated in simulation under diverse motor degradation and fault scenarios.
The rest of this paper is organized as follows. Section II describes the quadrotor dynamics and control framework. Section III presents the proposed motor efficiency estimator. Simulation results are given in Section IV, and conclusions are drawn in Section V.
II Modeling and Control of Quadrotor
II-A Quadrotor Dynamics
The quadrotor rigid-body dynamics are divided into translational and rotational components, capturing the effects of thrust, gravity, and inertia. The equations of motion are:
$\dot{x}$ $=v$ (1)
$m\dot{v}$ $=mge_{3}-f_{\text{c}}Re_{3}$ (2)
$\dot{R}$ $=R[\Omega]_{\times}$ (3)
$J\dot{\Omega}+\Omega\times J\Omega$ $=M,$ (4)
where $x\in\mathbb{R}^{3}$ is the position, $v\in\mathbb{R}^{3}$ is the velocity, $R\in\text{SO}(3)$ is the rotation matrix from the body-fixed frame to the inertial frame, $\Omega\in\mathbb{R}^{3}$ is the angular velocity, $M\in\mathbb{R}^{3}$ is the moment, $J\in\mathbb{R}^{3\times 3}$ is the inertia matrix, $f_{\text{c}}\in\mathbb{R}$ is the collective thrust, $g\in\mathbb{R}$ is the gravitational acceleration, and $e_{3}=[0,0,1]^{\top}$ is the unit vector along the inertial z-axis. The operator $\cdot$$_{\times}$ maps a vector to a skew-symmetric matrix $\mathbb{R}^{3}\mapsto\mathfrak{so}(3)$:
${}_{\times}\coloneqq\begin{bmatrix}0&-\Omega_{3}&\Omega_{2}\\ \Omega_{3}&0&-\Omega_{1}\\ -\Omega_{2}&\Omega_{1}&0\end{bmatrix}$. (5)
For simplicity, the inertia matrix in (4) is assumed diagonal:
$J=\operatorname{diag}(J_{xx},J_{yy},J_{zz}).$
II-B Geometric Tracking Control of Quadrotor
Following the geometric tracking control on $\text{SE}(3)$ developed by Lee et al. [17], a controller is designed to track desired trajectories. The tracking errors are defined as:
$e_{x}$ $=x-x_{d}$ (6)
$e_{v}$ $=v-v_{d}$ (7)
$e_{R}$ $=\frac{1}{2}(R_{d}^{\top}R-R^{\top}R_{d})^{\vee}$ (8)
$e_{\Omega}$ $=\Omega-R^{\top}R_{d}\Omega_{d},$ (9)
where $(\cdot)^{\vee}$ denote the inverse of (5), i.e., $\mathfrak{so}(3)\mapsto\mathbb{R}^{3}$, and $(x_{d},v_{d},R_{d},\Omega_{d})$ are the desired position, velocity, rotation matrix, and angular velocity. Based on (6)–(9), the control inputs, namely the collective thrust $f_{\text{c}}$ and moment $M$, are designed as:
$f_{\text{c}}=$ $-(-k_{x}e_{x}-k_{v}e_{v}-mge_{3}+m\ddot{x}_{d})\cdot Re_{3}$ (10)
$M=$ $-k_{R}e_{R}-k_{\Omega}e_{\Omega}+\Omega\times J\Omega$
$-J([\Omega]_{\times}R^{\top}R_{d}\Omega_{d}-R^{\top}R_{d}\dot{\Omega}_{d}),$ (11)
where $k_{x}$, $k_{v}$, $k_{R}$, $k_{\Omega}$, are control gains. The desired attitude $R_{d}$ is constructed by $R_{d}=[b_{2d}\times b_{3d};b_{2d};b_{3d}]\in\text{SO}(3)$, with the direction $b_{3d}$ to be designed as:
$b_{3d}=-\frac{-k_{x}e_{x}-k_{v}e_{v}-mge_{3}+m\ddot{x}_{d}}{\|-k_{x}e_{x}-k_{v}e_{v}-mge_{3}+m\ddot{x}_{d}\|}.$ (12)
Equation (12) defines the direction of the desired total thrust.
II-C Thrust Allocation and Motor Efficiency Modeling
The mapping between individual motor thrusts and the collective force and moments acting on the quadrotor can be represented by a linear transformation. Define the motor thrust vector:
$f_{\text{m}}=\begin{bmatrix}f_{1}&f_{2}&f_{3}&f_{4}\end{bmatrix}^{\top}$, (13)
where $f_{i}\in\mathbb{R}$ denotes the thrust by motor $i$ in (13). The actual collective thrust $f_{\text{c}}$ and moment $M=[M_{1},M_{2},M_{3}]^{\top}$ are related to the motor thrusts, scaled by an efficiency factor, with respect to the control inputs computed in (10)–(11):
$\begin{bmatrix}f_{\text{c},\text{actual}}\\ M_{1,\text{actual}}\\ M_{2,\text{actual}}\\ M_{3,\text{actual}}\end{bmatrix}=\Lambda Ef_{\text{m}}=\Lambda E\Lambda^{-1}\begin{bmatrix}f_{\text{c}}\\ M_{1}\\ M_{2}\\ M_{3}\end{bmatrix}$, (14)
where $\Lambda$ in (14) is the thrust allocation matrix, and $E$ is a diagonal matrix representing the motor efficiency. The thrust allocation matrix $\Lambda$ is defined as:
$\Lambda=\begin{bmatrix}1&1&1&1\\ -d&d&d&-d\\ d&d&-d&-d\\ -c_{\tau f}&c_{\tau f}&-c_{\tau f}&c_{\tau f}\end{bmatrix}$, (15)
where $d\in\mathbb{R}$ is the distance from the center of mass to each motor, and $c_{\tau f}\in\mathbb{R}$ is the thrust-to-torque coefficient in (15). The motor efficiency matrix $E$ is given by:
$E=\operatorname{diag}(\eta_{1},\eta_{2},\eta_{3},\eta_{4}).$ (16)
In (16), $\eta_{i}$ denotes the efficiency of motor ii. Assuming perfect efficiency ($\eta_{i}=1,\forall i$), the control commands $f_{\text{c}}$ and $M$ in (10)–(11) exactly match the actual $f_{\text{c, \text{actual}}}$ and $M_{\text{actual}}$ generated by the motors in (14).
II-D Numerical Integration of Quadrotor Dynamics
This section presents the numerical integration to predict states for motor efficiency estimation. The updates use the quadrotor dynamics model in (1)–(4), combined with the actual thrust and moment model computed in (14). The predicted states are later compared to measurements to construct trajectory residuals. Define the motor efficiency vector as:
$s=\begin{bmatrix}\eta_{1}&\eta_{2}&\eta_{3}&\eta_{4}\end{bmatrix}^{\top}$. (17)
The velocity is updated by integrating the translational acceleration over a small time interval $\Delta t\in\mathbb{R}$:
$\hat{v}_{t+1}(s)$ $=v_{t}+\int_{t}^{t+\Delta t}\bigg[ge_{3}-\frac{f_{\text{c},\text{actual}}Re_{3}}{m}\bigg]d\tau$
$\approx v_{t}+\bigg(ge_{3}-\frac{f_{\text{c},\text{actual}}Re_{3}}{m}\bigg)\Delta t.$ (18)
The position is updated by integrating the velocity and acceleration:
$\hat{x}_{t+1}(s)$ $=x_{t}+\int_{t}^{t+\Delta t}\bigg[v_{t}+\int_{t}^{\tau}ge_{3}-\frac{f_{\text{c},\text{actual}}Re_{3}}{m}d\tau^{\prime}\bigg]d\tau$
$\approx x_{t}+v_{t}\Delta t+\frac{1}{2}\bigg(ge_{3}-\frac{f_{\text{c},\text{actual}}Re_{3}}{m}\bigg)\Delta t^{2}.$ (19)
The angular velocity is integrated from the rotational dynamics:
$\hat{\Omega}_{t+1}(s)$ $=\Omega_{t}+\int_{t}^{t+\Delta t}\big[J^{-1}(M_{\text{actual}}-\Omega\times J\Omega)\big]d\tau$
$\approx\Omega_{t}+\big[J^{-1}(M_{\text{actual}}-\Omega\times J\Omega)\big]\Delta t.$ (20)
For rotation update, the integration of (3) is given by the matrix exponential:
$\hat{R}_{t+1}(s)=R_{t}\operatorname{exp}([\hat{\Omega}_{t}(s)]_{\times}\Delta t),$ (21)
where (21) can be expanded via a Taylor series:
$\text{exp}([\hat{\Omega}_{t}(s)]_{\times}\Delta t)=\sum_{k=0}^{\infty}\frac{([\hat{\Omega}_{t}(s)]_{\times}\Delta t)^{k}}{k!}.$ (22)
By neglecting higher order terms from (22), (21) can be approximated by:
$\hat{R}_{t+1}(s)\approx R_{t}(I+[\hat{\Omega}_{t}(s)]_{\times}\Delta t).$ (23)
III Motor Efficiency Estimation
III-A Overview of the Estimation Framework
Fig. 1 illustrates the proposed motor efficiency estimation framework. The estimator operates alongside a geometric tracking controller and takes as input the desired thrust and motor command computed by the controller. Within each sliding window of length $n$, the system collects the measured states including position $x$, velocity $v$, angular velocity $\Omega$, and rotation matrix $R$, together with the control inputs of collective thrust $f_{c}$, and moment $M$. These measurements are compared with model-based predictions, and the resulting residuals are minimized through constrained nonlinear optimization to estimate the efficiency of each motor.
Figure 1: Architecture of the proposed motor efficiency estimator integrated with geometric tracking controller.
III-B Definition of Trajectory Residuals
Trajectory residuals quantify the discrepancy between measured and predicted states from (18)–(20) and (23) with estimated efficiency vector $s$ in (17). The position, velocity, and angular velocity residuals are defined as:
$r_{v,t}(s)$ $=v_{t}-\hat{v}_{t}(s)$ (24)
$r_{x,t}(s)$ $=x_{t}-\hat{x}_{t}(s)$ (25)
$r_{\Omega,t}(s)$ $=\Omega_{t}-\hat{\Omega}_{t}(s).$ (26)
To define the orientation residual, the relative rotation between two consecutive time steps is extracted by:
$R_{t+1}$ $=R_{t}\delta R$
$\rightarrow\delta R$ $=R_{t}^{\top}R_{t+1}.$ (27)
On the other hand, assuming small $\Delta t$, the predicted incremental rotation is approximated by:
$\delta\hat{R}(s)$ $=\operatorname{exp}([\hat{\Omega}_{t}(s)]_{\times}\Delta t)$
$\approx I+[\hat{\Omega}_{t}(s)]_{\times}\Delta t.$ (28)
With (27) and (28), the orientation residual is defined as:
$r_{R,t}(s)$ $=\frac{1}{2}\operatorname{tr}[I-\delta R^{\top}\delta\hat{R}(s)].$ (29)
Stacking (24)–(26) and (29) over a sliding window of $n$-segments gives:
$r(s)=\frac{1}{\sqrt{n}}\begin{bmatrix}r_{v,t=0}(s)\\ r_{x,t=0}(s)\\ r_{\Omega,t=0}(s)\\ r_{R,t=0}(s)\\ \vdots\\ r_{v,t=n-1}(s)\\ r_{x,t=n-1}(s)\\ r_{\Omega,t=n-1}(s)\\ r_{R,t=n-1}(s)\end{bmatrix}$. (30)
The aggregated residual vector in (30) is used to estimate motor efficiency factors via optimization.
III-C Formulation of the Optimization Problem
The motor efficiency estimation is formulated as a constrained nonlinear optimization problem. The objective function comprises two terms: (i) a trajectory residual penalty term and (ii) a temporal smoothness regularization term that promotes temporal smoothness across successive estimates. The cost function at time $t$ is defined as:
$F(s_{t})=\frac{1}{2}\big\|r(s_{t})\big\|^{2}_{G}+\frac{\gamma}{2}\big\|s_{t}-s_{t-1}\big\|^{2},$ (31)
where $\|r(s_{t})\|_{G}=\sqrt{r(s_{t})^{\top}Gr(s_{t})}$ denotes the weighted norm of the residual vector $r(s_{t})$, and $G\in\mathbb{R}^{10n\times 10n}$ is a positive semidefinite block-diagonal weighting matrix. The local weight matrix for each time segment is:
$G_{i}=\operatorname{diag}($ $g_{v_{1}},g_{v_{2}},g_{v_{3}},g_{x_{1}},g_{x_{2}},g_{x_{3}},$
$g_{\Omega_{1}},g_{\Omega_{2}},g_{\Omega_{3}},g_{R}).$ (32)
Using (32), the full weighting matrix is assembled as:
$G=\operatorname{blockdiag}(w_{0}G_{0},\cdots,w_{n}G_{n}),$ (33)
where $w_{i}$ is the weight for outlier rejection that will be introduced in the next section. Finally, with the objective function (31) with $G$-norm defined in (33), the optimization problem is defined as:
$\min_{s_{t}}$ $F(s_{t})$ (34)
s.t. $\eta_{\text{min}}\leq\eta_{i}\leq\eta_{\text{max}},$
$i=1,\cdots,4.$
These constraints ensure that each motor’s efficiency remains within physically valid bounds.
III-D Outlier Rejection and Reweighting
To suppress the influence of abnormal measurements, a median absolute deviation (MAD)–based reweighting scheme is applied at each outer iteration. For the $i$-th segment, the weighted residual energy is
$e_{i}=r_{i}(s)^{\top}G_{i}r_{i}(s).$ (35)
The residual energy median and MAD are computed as
$\tilde{m}$ $=\operatorname{median}(e_{i}),$ (36)
[…]
Each residual energy is normalized into a robust z-score:
$z_{i}=\frac{|e_{i}-\tilde{m}|}{\max(\operatorname{\text{MAD}},\epsilon_{\text{min}})}$ (37)
and the corresponding weight is updated by
$w_{i}=\max\bigg(\frac{1}{1+(\frac{z_{i}}{z_{\text{soft}}})^{p}},w_{\text{min}}\bigg),$ (38)
where $z_{\text{soft}}$ controls the decay threshold, $p$ specifies the decay rate, and $w_{\text{min}}$ denotes the minimum weight. To further discard extreme outliers, a hard rejection is imposed:
$w_{i}=\begin{cases}0&\text{if\ }z_{i}>z_{\text{hard}}\\ w_{i}&\text{otherwise}.\end{cases}$ (39)
(35)-(39) jointly apply soft decay and hard rejection, which attenuate moderate outliers while eliminating extreme ones, thereby improving robustness to sensor noise and modeling errors.
III-E KKT Conditions for Constrained Optimization
To solve the constrained optimization problem of (34), a primal-dual interior-point method is adopted. The inequality constraints are reformulated in the standard form:
$\phi(s_{t})\preceq 0,$ (40)
where (40) is defined as:
$\phi(s)=\begin{bmatrix}\phi_{1}(s)\\ \phi_{2}(s)\\ \phi_{3}(s)\\ \phi_{4}(s)\\ \phi_{5}(s)\\ \phi_{6}(s)\\ \phi_{7}(s)\\ \phi_{8}(s)\\ \end{bmatrix}=\begin{bmatrix}\eta_{1}-\eta_{\text{max}}\\ \eta_{2}-\eta_{\text{max}}\\ \eta_{3}-\eta_{\text{max}}\\ \eta_{4}-\eta_{\text{max}}\\ \eta_{\text{min}}-\eta_{1}\\ \eta_{\text{min}}-\eta_{2}\\ \eta_{\text{min}}-\eta_{3}\\ \eta_{\text{min}}-\eta_{4}\end{bmatrix}$. (41)
A corresponding dual variable $\lambda\in\mathbb{R}^{m}$, $m=8$ is introduced, associated with the inequality constraints defined by (40) and (41). The Karush–Kuhn–Tucker (KKT) conditions for optimality consist of the following:
$\bullet$ Stationarity:
$\nabla F(s_{t})+\sum_{i=1}^{m}\lambda_{i}\nabla\phi_{i}(s_{t})=0.$ (42)
$\bullet$ Primal feasibility:
$\phi_{i}(s_{t})$ $\leq 0,\ i=1,\cdots,m.$ (43)
$\bullet$ Dual feasibility:
$\lambda_{i}$ $\geq 0,\ i=1,\cdots,m.$ (44)
$\bullet$ Complementary slackness:
$\lambda_{i}$ $\phi_{i}(s_{t})=0,\ i=1,\cdots,m.$ (45)
The conditions of (42)–(45) are enforced iteratively using a primal-dual interior-point method, which ensures convergence to a solution satisfying optimality and feasibility. Details of the solution strategy are provided in the following section.
III-F Primal-Dual Interior-Point Method
To handle the inequality-constrained optimization in (34), the problem is reformulated using a logarithmic barrier function:
$\min_{s_{t}}\quad F(s_{t})-\frac{1}{\beta}\sum_{i=1}^{8}\text{log}\big(-\phi_{i}(s_{t})\big).$ (46)
The stationarity condition of (46) is derived as:
$\nabla F(s_{t})+\frac{1}{\beta}\sum_{i=1}^{8}\bigg(\frac{1}{-\phi_{i}(s_{t})}\nabla\phi_{i}(s_{t})\bigg)=0.$ (47)
By observing (47), the dual variables $\lambda_{i}$ can be expressed as:
$\lambda_{i}=\frac{1}{-\beta\phi_{i}(s_{t})}.$ (48)
Rearranging (48) yields:
$\lambda_{i}\phi_{i}(s_{t})=-\frac{1}{\beta}$ (49)
$\rightarrow$ $-\lambda_{i}\phi_{i}(s_{t})-\frac{1}{\beta}=0.$
Equation (49) is called the modified complementary slackness condition, which approaches the standard complementary slackness condition (45) as $\beta\rightarrow\infty$, forming the central path followed by the primal-dual interior-point method. The KKT residual vector is defined as:
$r_{\text{KKT}}(s,\lambda)=\begin{bmatrix}\nabla F(s_{t})+D\phi(s_{t})^{\top}\lambda\\ -\operatorname{diag}(\lambda)\phi(s_{t})-(1/\beta)\mathbf{1}\end{bmatrix}$
$=\begin{bmatrix}J_{r}(s_{t})^{\top}Gr(s_{t})+\gamma(s_{t}-s_{t-1})+D\phi(s_{t})^{\top}\lambda\\ -\operatorname{diag}(\lambda)\phi(s_{t})-(1/\beta)\mathbf{1}\end{bmatrix}$, (50)
where $D\phi(s)$ is the Jacobian matrix of constraints in (50):
$D\phi(s)=\begin{bmatrix}\nabla\phi_{1}(s)^{\top}\\ \nabla\phi_{2}(s)^{\top}\\ \nabla\phi_{3}(s)^{\top}\\ \nabla\phi_{4}(s)^{\top}\\ \nabla\phi_{5}(s)^{\top}\\ \nabla\phi_{6}(s)^{\top}\\ \nabla\phi_{7}(s)^{\top}\\ \nabla\phi_{8}(s)^{\top}\\ \end{bmatrix}=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\\ -1&0&0&0\\ 0&-1&0&0\\ 0&0&-1&0\\ 0&0&0&-1\end{bmatrix}$. (51)
Along with (51), the Jacobian of trajectory residuals $J_{r}(s)$ in (50) is constructed by stacking residual derivatives across $n$ segments:
$J_{r}(s)=\begin{bmatrix}J_{r,0}(s)\\ \vdots\\ J_{r,n-1}(s)\end{bmatrix},\ J_{r,i}(s)=\begin{bmatrix}\frac{\partial r_{v,t}(s)}{\partial s}\\ \frac{\partial r_{x,t}(s)}{\partial s}\\ \frac{\partial r_{\Omega,t}(s)}{\partial s}\\ \frac{\partial r_{R,t}(s)}{\partial s}\end{bmatrix}$. (52)
Defining primal-dual variable $y=[s,\lambda]^{\top}$ and apply Newton’s method to linearize the KKT residuals:
$r_{\text{KKT}}(y+\Delta y)\approx r_{\text{KKT}}(y)+Dr_{\text{KKT}}(y)\Delta y=0,$ (53)
which leads to the following linear system:
$\underbrace{\begin{bmatrix}J_{r}(s_{t})^{\top}GJ_{r}(s_{t})+\gamma I&D\phi(s_{t})^{\top}\\ -\operatorname{diag}(\lambda)D\phi(s_{t})&-\operatorname{diag}(\phi(s_{t}))\end{bmatrix}}_{Dr_{\text{KKT}}(y)}\underbrace{\begin{bmatrix}\Delta s_{t}\\ \Delta\lambda\end{bmatrix}}_{\Delta y}=-\underbrace{\begin{bmatrix}r_{\text{dual}}\\ r_{\text{cent}}\end{bmatrix}}_{r_{\text{KKT}}(y)}.$ (54)
Solving the Newton system (53)–(54) gives the primal-dual update direction:
$\Delta y=-(Dr_{\text{KKT}}(y))^{-1}r_{\text{KKT}}(y).$ (55)
To preserve feasibility ($\lambda_{i}\geq 0$), the maximum step size $\alpha_{\text{max}}$ for (55) is determined as:
$\alpha_{\text{max}}$ $=\operatorname{sup}\{\alpha\in[0,1]\ |\ \lambda+\alpha\Delta\lambda\succeq 0\}$
$=\min\{1,\min\{-\lambda_{i}/\Delta\lambda_{i}|\Delta\lambda_{i}<0\}\}.$ (56)
With (56), a backtracking line search begins with $\alpha=0.99\alpha_{\text{max}}$ and iteratively reduces $\alpha$ by a factor $\zeta\in(0,1)$ until the following condition is met:
$\big\|r_{\text{KKT}}(s_{t}^{+},\lambda^{+})\|\leq(1-\kappa\alpha)\|r_{\text{KKT}}(s_{t},\lambda)\|+\epsilon_{\text{tol}},$ (57)
where $\kappa\in(0,1)$ and $\epsilon_{\text{tol}}>0$ is a tolerance parameter accounting for noise in the data. The updates used in (57) are:
$s_{t}^{+}=s_{t}+\alpha\Delta s_{t},\ \lambda^{+}=\lambda+\alpha\Delta\lambda.$
The surrogate duality gap is computed as:
$\hat{\delta}(s_{t},\lambda)=-\phi(s_{t})^{\top}\lambda.$ (58)
Equation (58) serves as a convergence indicator and as $\hat{\delta}(s_{t},\lambda)$ decreases towards zero, the iterates approach primal and dual feasibility. The optimization terminates when $\|r_{\text{dual}}\|$ and $\hat{\delta}(s_{t},\lambda)$ falls below specified tolerances. The overall primal-dual interior-point method is summarized in Algorithm 1.
Input: Feasible $s_{t}$, $\lambda>0$, $\mu>1$, $\epsilon_{\text{feas}}>0$, $\epsilon_{\text{gap}}>0$
Output: Optimized $s_{t}$ and $\lambda$
1 repeat
2   Set barrier parameter: $\beta\leftarrow\mu m/\hat{\delta}(s_{t},\lambda)$
3   Compute the primal-dual direction: $\Delta y$
4   Determine step size $\alpha$ via line search
5   Update variables: $y\leftarrow y+\alpha\Delta y$
6
7until $\|r_{\text{dual}}\|\leq\epsilon_{\text{feas}}$ and $\hat{\delta}(s_{t},\lambda)\leq\epsilon_{\text{gap}}$
Algorithm 1 Primal-Dual Interior-Point Method
III-G Iteratively Reweighted Least Squares (IRLS)
To enhance robustness, the interior-point method is embedded in an outer IRLS loop. At each iteration, residuals are scored with a MAD-based robust z-score to down-weight outliers and discard extreme segments. The reweighted problem is then solved by the inner primal-dual interior-point method (i.e., Algorithm 1), with convergence achieved in about three outer iterations. The overall procedures of IRLS are presented in Algorithm 2.
Input: Feasible $s^{(0)}_{t}$, Max iterations $N_{\text{IRLS}}$
Output: Optimized $s_{t}$
1 for $k=1$ to $N_{\text{IRLS}}$ do
2   Compute trajectory residuals $r_{i}(s^{(k-1)}_{t})$
3   Compute weighted residual energy $e_{i}$
4   Compute robust z-score $z_{i}$
5   Update weights $w_{i}$ and hard-reject if $z_{i}>z_{\text{hard}}$
6   Form full block-diagonal weight matrix $G$
7   Run primal-dual interior-point method for $s^{(k)}_{t}$
8
return $s^{(k)}_{t}$
Algorithm 2 Iteratively Reweighted Least Squares
III-H Derivatives of Trajectory Residuals
This section derives the partial derivatives of all residuals with respect to the motor efficiency parameters $s$. The derivative of the velocity residual with respect to $s$ is:
$\frac{\partial r_{v,t}(s)}{\partial s}=\frac{\Delta t}{m}\begin{bmatrix}f_{1}r_{13}&f_{2}r_{13}&f_{3}r_{13}&f_{4}r_{13}\\ f_{1}r_{23}&f_{2}r_{23}&f_{3}r_{23}&f_{4}r_{23}\\ f_{1}r_{33}&f_{2}r_{33}&f_{3}r_{33}&f_{4}r_{33}\\ \end{bmatrix}$. (59)
The derivative of the position residual with respect to $s$ is:
$\frac{\partial r_{x,t}(s)}{\partial s}=\frac{\Delta t^{2}}{2m}\begin{bmatrix}f_{1}r_{13}&f_{2}r_{13}&f_{3}r_{13}&f_{4}r_{13}\\ f_{1}r_{23}&f_{2}r_{23}&f_{3}r_{23}&f_{4}r_{23}\\ f_{1}r_{33}&f_{2}r_{33}&f_{3}r_{33}&f_{4}r_{33}\\ \end{bmatrix}$. (60)
The derivative of the angular velocity residual with respect to $s$ is:
$\frac{\partial r_{\Omega,t}(s)}{\partial s}=\Delta t\begin{bmatrix}\frac{f_{1}d}{J_{xx}}&\frac{-f_{2}d}{J_{xx}}&\frac{-f_{3}d}{J_{xx}}&\frac{f_{4}d}{J_{xx}}\\ \frac{-f_{1}d}{J_{yy}}&\frac{-f_{2}d}{J_{yy}}&\frac{f_{3}d}{J_{yy}}&\frac{f_{4}d}{J_{yy}}\\ \frac{f_{1}c_{\tau f}}{J_{zz}}&\frac{-f_{2}c_{\tau f}}{J_{zz}}&\frac{f_{3}c_{\tau f}}{J_{zz}}&\frac{-f_{4}c_{\tau f}}{J_{zz}}\\ \end{bmatrix}$. (61)
Recall (27) and define the rotation matrix difference:
$\delta R^{\top}\coloneqq\begin{bmatrix}\delta r_{11}&\delta r_{21}&\delta r_{31}\\ \delta r_{12}&\delta r_{22}&\delta r_{32}\\ \delta r_{13}&\delta r_{23}&\delta r_{33}\\ \end{bmatrix}$. (62)
Based on (62), the derivative of the rotation residual with respect to $s$ is:
$\frac{\partial r_{R,t}(s)}{\partial s}=$ (63)
$\frac{\Delta t^{2}}{2}\begin{bmatrix}-\frac{(\delta r_{23}-\delta r_{32})f_{1}d}{J_{xx}}+\frac{(\delta r_{31}-\delta r_{13})f_{1}d}{J_{yy}}-\frac{(\delta r_{12}-\delta r_{21})f_{1}c_{\tau f}}{J_{zz}}\\ \frac{(\delta r_{23}-\delta r_{32})f_{2}d}{J_{xx}}+\frac{(\delta r_{31}-\delta r_{13})f_{2}d}{J_{yy}}+\frac{(\delta r_{12}-\delta r_{21})f_{2}c_{\tau f}}{J_{zz}}\\ \frac{(\delta r_{23}-\delta r_{32})f_{3}d}{J_{xx}}-\frac{(\delta r_{31}-\delta r_{13})f_{3}d}{J_{yy}}-\frac{(\delta r_{12}-\delta r_{21})f_{3}c_{\tau f}}{J_{zz}}\\ -\frac{(\delta r_{23}-\delta r_{32})f_{4}d}{J_{xx}}-\frac{(\delta r_{31}-\delta r_{13})f_{4}d}{J_{yy}}+\frac{(\delta r_{12}-\delta r_{21})f_{4}c_{\tau f}}{J_{zz}}\end{bmatrix}$.
With the analytical expressions derived in (59)–(61) and (63), the Jacobian $J_{r}(s)$ in (52) can be constructed to compute the Newton step in the primal-dual interior-point method.
III-I EKF Baseline
For comparison, we implement an extended Kalman filter (EKF) as baseline. The EKF state is defined as
$s_{\mathrm{EKF}}=\begin{bmatrix}x^{\top}&v^{\top}&\Omega^{\top}&\mathrm{vec}(R)^{\top}&\eta^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{22},$
where the dynamics of motor efficiency are modeled as random walks:
$\dot{\eta}=w_{\eta},\quad w_{\eta}\sim\mathcal{N}(0,Q_{\eta}).$
This EKF serves as a filter-based baseline for comparison against the proposed batch optimization approach.
IV Simulations
This section presents simulation results to validate the effectiveness of the proposed motor efficiency estimation framework under various degradation and fault scenarios. The MATLAB code used for the simulations can be accessed at https://github.com/shengwen-tw/irls-quadrotor-motor-efficiency-estimator.
IV-A Simulation Setup
The quadrotor parameters and controller gains are configured to resemble the characteristics of a typical F450 platform, as summarized in Table I.
TABLE I: Quadrotor Parameters and Control Gains
Parameter Value
$m$ 1.0 kg
$J$ $\operatorname{diag}(0.01466,0.01466,0.02848)\ \text{kg}\cdot\text{m}^{2}$
$d$ 0.225 m
$c_{\tau f}$ 0.009012m0.009012\ $\text${m}
$k_{x}$ $\operatorname{diag}(9,9,12)$
$k_{v}$ $\operatorname{diag}(7,7,12)$
$k_{R}$ $\operatorname{diag}(10,10,10)$
$k_{\Omega}$ $\operatorname{diag}(2,2,2)$
The desired trajectory consists of a circular motion in the $xy$-plane combined with a constant value in the $z$-direction, defined as:
$x_{d}(t)=\begin{bmatrix}3\cos(0.2\pi t)\\ 3\sin(0.2\pi t)\\ -1\end{bmatrix},\ v_{d}(t)=\begin{bmatrix}-0.6\pi\sin(0.2\pi t)\\ 0.6\pi\cos(0.2\pi t)\\ 0\end{bmatrix}$.
The desired body x-axis direction and angular velocity are given by:
$b_{1d}(t)=\begin{bmatrix}\cos(0.1\pi t)\\ \sin(0.1\pi t)\\ 0\end{bmatrix},\ \Omega_{d}(t)=\begin{bmatrix}0\\ 0\\ 0\end{bmatrix}$.
To evaluate the robustness of the proposed estimator, the following conditions are simulated:
$\bullet$ Voltage-Induced Gradual Degradation
Motor efficiency $\eta_{i}(t)$ degrades exponentially with battery voltage $V(t)\in\mathbb{R}$:
$\eta_{i}(t)=\eta_{i}(0)\exp(\xi(V(t)-V(0))),\ i=1,\cdots,4,$
where $\eta_{i}(0)$ is the initial efficiency and $\xi>0$ is a voltage sensitivity constant.
$\bullet$ Abrupt Fault Injection
At specified time intervals, selected motors experience sudden efficiency drops:
[…]
$\bullet$ Random Thrust Noise
To emulate real-world disturbances, random noise is added to the motor thrust vector:
$f_{i,\epsilon_{f}}=f_{i}\cdot\exp(\epsilon_{f}),\ \epsilon_{f}\sim\mathcal{N}(0,\sigma_{f}),$
where $\sigma_{f}$ denotes the thrust noise level.
IV-B Simulation Results and Transition Robustness
Fig. 2-4 compare the EKF baseline and the proposed estimator under gradual degradation, abrupt fault injection, and combined cases, all with thrust noise level $\sigma_{f}=0.07$. While both methods achieve similar long-term accuracy when calibrated to the same root mean square error (RMSE) and standard deviation levels, their behavior during transitions differs markedly. The EKF exhibits pronounced spikes whenever abrupt changes occur, as incoherent measurements are directly incorporated into the filter updates. In contrast, the proposed IRLS-based estimator effectively down-weights outliers and maintains smooth estimates through transitions, accurately reflecting the true motor efficiency even under sudden faults.
IV-C RMSE and Spike Comparison Across Scenarios
Fig. 5 summarizes the RMSE and standard deviation across multiple scenarios, with both methods calibrated to the same accuracy level. Under this normalization, Fig. 6 highlights the main difference: the proposed method consistently suppresses the transient spikes observed in the EKF during fault injection and combined scenarios. These results confirm that the proposed approach offers robustness advantages while preserving comparable average accuracy.
IV-D Estimation and Optimization Convergence
Fig. 7 and Fig. 8 illustrate the convergence properties of the proposed estimator. Fig. 7 shows that starting from an initial guess of $0.5$, the efficiency estimates converge to the true values within one full loop of the interior-point method. Fig. 8 further presents the evolution of primal residual, dual residual, and surrogate duality gap, all of which decrease monotonically within the same loop. Together, these results demonstrate that the optimization reliably satisfies KKT optimality conditions and converges efficiently.
Figure 2: Comparison between the EKF baseline and the proposed estimator under gradual voltage-induced degradation with thrust noise.
Figure 3: Comparison between the EKF baseline and the proposed estimator under abrupt motor fault injection with thrust noise.
Figure 4: Comparison between the EKF baseline and the proposed estimator under combined degradation and abrupt fault scenarios with thrust noise.
Figure 5: Root mean square error (RMSE) and standard deviation of motor efficiency estimation under various degradation and fault scenarios, comparing the EKF baseline with the proposed method.
Figure 6: Comparison of maximum estimation spikes for each motor between the EKF baseline and the proposed method under (a) abrupt-only faults and (b) combined degradation-and-abrupt faults.
Figure 7: Convergence of motor efficiency estimates from an initial guess of $0.5$ over one full loop of the interior-point method.
Figure 8: Evolution of primal residual, dual residual, and surrogate duality gap over one full loop of the interior-point method, confirming KKT convergence.
V Conclusion
This paper presented a data-driven framework for online estimation of quadrotor motor efficiency based on residual minimization. The estimation problem was formulated as a constrained nonlinear optimization within a sliding-window structure, solved efficiently using a primal-dual interior-point method. To enhance robustness, an outer iteratively reweighted least squares (IRLS) loop with robust z-score weighting was integrated, enabling rejection of incoherent measurements and suppression of transient spikes.
Compared to traditional filter-based methods such as EKF, the proposed batch-mode optimization achieves similar long-term accuracy while exhibiting markedly improved robustness during abrupt transitions. Simulation results under degradation, fault injection, and combined scenarios showed that the estimator avoids the large spikes observed in EKF, particularly in motor clipping cases, by effectively rejecting outliers.
The methodology is well-suited for applications in fault detection and isolation, health monitoring, and predictive maintenance in aerial robotic systems. Its formulation supports extensibility toward learning-enabled estimation and control architectures. Future research may focus on enhancing robustness against unmodeled effects such as wind disturbances and external perturbations.